The income for a company producing electronic components is given by
y= -20x2-50x+200, where x is the price in dollars of each component. The cost for the production is given by y = 60x −10 . Determine the price that will allow the production to break even.

Answer by Boreal(15235)

Set the two equal to one another.
-20x^2-50x+200=60x-10
-20x^2-110x+210=0
2x^2+11x-21=0
(2x-3)(x+7)=0
positive root is 3/2 of $1.50